# A Review On Safe Reinforcement Learning Using Lyapunov and Barrier Functions

### v2

@@ Abstract
{IEEEImpStatement}
Reinforcement learning and data-driven control has seen significant research over the last two decades given its usability and effectiveness for computing control policies in unknown stochastic environments. Given most real-world scenarios are difficult to model and further have significant unknown disturbances, the problem becomes especially challenging. Safe reinforcement learning provides an elegant solution to these challenges as agents can safely explore and tackle these environments which creates several use cases for real-world applications as covered in this review.
Reinforcement learning in its bare-bones fails to satisfy or provide formal guarantees for safety and stability with respect to the computed policy. Lyapunov and barrier functions have long been employed in control theory and optimization problems to guarantee system stability and satisfy state constraints, respectively. Incorporating Lyapunov and barrier methods in reinforcement learning helps to address the issues of safety and stability to increase and promote its use in safety critical real-world applications.
The advantages of safe reinforcement learning extend directly to real-world applications and provide exciting opportunities to deploy agents which can learn safely in real-time, as discussed and covered in this review. Most of the literature in this field is fairly recent, dating back to 2017. This review compiles and presents important theoretical developments and challenges in the field and, further proposes possible directions for improving existing formulations.
{IEEEkeywords}
Reinforcement Learning, Lyapunov Functions, Barrier Functions, Neural Networks (NNs), Artificial Intelligence Safety, Autonomous agents, Intelligent control.
@@ 1 Introduction
With increase in efficiency and function approximation capabilities of deep neural networks (DNN), reinforcement learning has seen increased research and some exciting developments over the last few decades [1]. RL in its bare bones involves an agent iteratively interacting with an environment to compute a control or decision making policy by maximizing a reward function. RL has proven to be effective in computing control or decision policies for a wide variety of complex systems and environments such as robotics [2, 3, 4, 5, 6], computer vision [7, 8, 9], cyber security [10, 11, 12, 13], energy management [14, 15, 16, 17]), chess [18, 19, 20] and video games [21, 22, 23] to name a few. However, when using RL there are several complications and theoretical gaps in terms of reproducibility, convergence guarantees, large amounts of data for training and large number of iterations required for convergence [24]. Even though RL is often known for its data-driven nature, it can be broadly classified into two approaches for deriving a control or decision policy namely, model-based (value and policy iteration) and model-free (value-function, policy search or gradient and actor-critic methods) approaches. Model-based approaches imply use of knowledge (partial or complete) or an approximation of the system to derive a control policy whereas, model-free approaches purely rely on data or experiences collected by interactions with the environment without using a model of the system [25]. It seems intuitive to have stronger guarantees for constraint satisfaction and computing a stabilizing policy while leveraging a model-based approach as the RL agent uses information from the system model. There has been some success in providing these guarantees for model-free approaches as well, notably in [26]. A key challenge for policies derived through use of RL has been providing guarantees for stability of the system for a computed policy and satisfying system state constraints. Broadly speaking this challenge encapsulates the need for safety in RL for real-world applications and thus, encourages the research in this field.
@@
There are several methods in literature for computing a stabilizing policy and satisfying system constraints in RL. The simplest approaches offer a direct and intuitive method of filtering out “unsafe” actions and policies. The term “unsafe” can be referred to any action leading to constraint violation or system instability. In [27], the authors propose a confidence safety filter for stochastic nonlinear systems. The key idea presented is to formulate safety constraints as cumulative costs which can then be expressed as cost constraints for RL. Backup policies in case of constraint violations are obtained for the safety filter by computing a safe policy through robust RL, which is then employed in a confidence-based safety filter. The authors in [28], use a convex optimization-based filtering to satisfy hard constraints for target tracking. They propose a two-level motion planner where a RL-based controller generates an input based on the target and quadratic programming (QP) based safety filter generates a safe input, regulation is performed on these two inputs produced. Similar to safety filter, supervisor-based safe RL methods often use human supervision [29, 30, 31] to penalize or replace unsafe actions with desirable ones. These supervisor-based methods are often termed as human-in-loop RL. A formal approach to supervisor-based safe RL methods are classified as RL via shielding. Authors in [32], propose a centralized (single shield for all agents in an environment) and factored (separate shields for a subsets of agents in an environment) shielding methodology for a multi-agent RL setting using linear temporal logic (LTL) for safety constraints. The scalability and learning performance are discussed for each methodology by the authors of mentioned work. In [33], authors express constraints as temporal logic and propose an algorithm for automatic synthesis of shields for the given temporal logic. This is further incorporated with RL and the results show increased learning performance compared to unshielded cases. However, a key drawback of this approach is that it relies on some prior information for model of the system to deem an action unsafe. Model predictive control (MPC) has long been effective in control theory for handling uncertainty in dynamical systems while satisfying operational constraints. There have been several notable works in combining the data-driven advantage of RL and the robustness of MPC, reviewed in [34]. RL using robust MPC is proposed in [35], where the authors use online data to compute safe design constraints to evaluate the cost function for RL, Q-function and value function are obtained via robust MPC. In [36], combination of QP and MPC, one at a time with RL is explored. The policy and value functions of RL are approximated using MPC and QP instead of employing DNNs. Compared to lack of guarantees and explainability of DNNs, this approach provides structure to the approximations for policy and value functions.
@@
Control Lyapunov functions (CLF) [37] and control barrier functions (CBF) [38] have proven to be an effective way to guarantee safety through stability of a closed-loop system and defining safe sets, respectively. Ensuring safe operation for dynamical systems involving human interaction, expensive equipment etc., requires strong guarantees during and after the learning process for RL agents. Using CBFs and CLFs to provide stability and safety guarantees is especially well suited for learning because both features may be stated using learnable functions. Considering their usefulness in a learning framework and their applicability to a wide variety of control frameworks, this paper reviews methods using Lyapunov and barrier functions to ensure safety and stability guarantees of a computed policy while using RL for decision-making problems. Furthermore, most real-world applications require safety during the training process (for example, navigation in unknown environments for drones, human presence in application etc.) which makes this problem crucial as it provides theoretical guarantees for safety. The key motivation for this review is to discuss methods employed in recent work, not covered in reviews and surveys mentioned in Section 2 and particularly focus on use of Lyapunov and barrier functions for RL to guarantee satisfying system constraints and stability in terms of the computed policy. We focus more on the theoretical approaches present in literature rather than the type of applications it is employed. To the best of author’s knowledge, such a review focusing on use of Lyapunov and barrier functions in safe RL does not exist and this research gap is addressed in this review. Rest of the review is organized as follows: Section 2 presents scope of work for different reviews and surveys considered. Section 3 discusses preliminary theory and notations briefly to provide the reader a background for further discussion. Section 4 presents and summarizes the current and prior work related to use of: (i) Lyapunov functions in RL; (ii) Barrier functions in RL; (iii) Lyapunov and barrier functions in RL. Section 5 provides a discussion on future research directions, current challenges and potential real-world applications. Finally, Section 6 presents conclusions to the review conducted.
@@ 2 Related Work
The literature in safe reinforcement learning has been fairly recent, spanning back to the last two decades. The earliest survey in this field by [39], primarily focuses on safe RL by augmenting the optimization criterion and exploration process by including a risk parameter and state constraints. The survey also focuses on leveraging external knowledge in the form of expert policies and human intervention to guide the exploration process. The review by Liu et al. [40] focuses on policy learning with constraints for model-free RL. The authors present a robust taxonomy for the types of constraints for a constrained Markov decision process (CMDP) and discusses various policy optimization approaches based on the type of constraints in CMDP. The authors broadly classifies constraints as cumulative and instantaneous depending on the long-term and short-term effects of constraints on the cost function, respectively.
@@
Brunke et al. [41] provides an extensive and one of the most comprehensive reviews covering safe learning for control from a robotics perspective, however, they do not cover key recent and prior works on combining Lyapunov and barrier functions in RL. To quote the authors, they provide a “bird’s eye view” of the field. Authors cover most of the literature associated with safe learning and provide excellent categorizations for levels of safety and various formulations for safe learning techniques. However, due to the vast range of topics considered, use of Lyapunov and barrier functions in RL is sparsely covered. The review does provide an excellent starting point for an overview into all possible directions for development in safe learning for control and RL. Gu et al. [42] focus on all developments in safe RL covering theory and applications. The authors formalize a “2H3W” problem, addressing the key problems and challenges pertaining to safe RL implementation. Convergence analysis and iteration complexity for various approaches to safe RL are explored for model-free and model-based RL along with applications and available benchmarks. The authors also discuss application of safe RL in multi-agent settings, discussing possible research avenues. The review is detailed, however, it covers only a few approaches pertaining to use of Lyapunov and barrier functions.
@@
The authors in [25], present an overview for RL with guarantees. The authors cover three approaches for stability of the system for a computed policy and constraint satisfaction namely, RL with supervisors, model predictive control (MPC)-based RL and Lyapunov-based RL. The review presents few instances of using CBF with CLF, but it primarily gives an overview of key challenges in consideration of model-based and model-free RL while providing safety guarantees. Anand et al. [43] presents a concise and focused review on safe learning for control using CLF and CBF. The authors cover safety in RL, online supervised learning (SL) and offline SL with incorporation of CLF, CBF and combining CLF with CBF. However, the review presents a few notable approaches for combining CLF and CBF with RL and only covers one CMDP based formulation. The excellent survey by Dawson et al. [44] covers a great depth of literature available on learning-based control using CLF, CBF and contraction methods. The review provides excellent theoretical overview of different formulations and implementations using control certificates but is more focused towards learning-based control (adaptive controls) from a control theoretic perspective compared to RL.
@@ 3.1 Definition of Safety
A general continuous time nonlinear dynamical system is represented as $\dot{x}=f(x,u)$, where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is state of the system, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control input and $f:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ is the state transition function (assumption: Locally Lipschitz in $x$ and $u$) [44]. The control affine form of nonlinear dynamical system can represented as $\dot{x}=f(x)+g(x)u$, where $g:\mathbb{R}^{n}\mapsto\mathbb{R}^{n\times m}$ (assumption: Locally Lipschitz in $x$ and $u$).
@@ 3.2 Constrained Markov Decision Process (CMDP) and Safe RL
A Markov Decision Process (MDP) can be denoted by the tuple $\mathcal{S},\mathcal{A},\mathcal{R},\mathbb{P},\mu,\gamma>$  [45], where $\mathcal{S}$ and $\mathcal{A}$ denote the set of states and actions, respectively. $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ denotes the reward function, $\mathbb{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ denotes the transition probability function, $\mu:\mathcal{S}\mapsto[0,1]$ is the initial probability distribution and $\gamma$ denotes the discount factor for future rewards. A policy $\pi:\mathcal{S}\mapsto\mathcal{P}(A)$ is a mapping from states to a probability distribution over actions and $\pi(a_{t}|s_{t})$ is the probability of taking action $a$ under state $s$ at time $t$. Reinforcement learning (RL) involves an agent successively interacting with its environment, to accomplish the aim for learning a policy to maximize a given reward criteria. The goal of RL is to learn a policy $\pi$ that maximizes the discounted cumulative reward:
@@
where $\tau$ denotes a state-action trajectory, and $\tau\sim\pi$ denotes trajectories sampled from the policy $\pi$. A CMDP is an extension of MDP with an additional constraint set of cost functions $\mathcal{C}$ [46]. The augmented tuple for a CMDP is then denoted by $\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{C},\mathbb{P},\mu,\gamma>$. The cost functions $\mathcal{C}_{i}\in\mathcal{C}$, $\mathcal{C}_{i}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ are constraint associated cost functions. We call an action $a$ feasible if $a\in\mathcal{A}$ satisfies all of its constraints. The objective for safe RL is then to find a policy to maximize the long-term rewards while satisfying the constraints, denoted as:
@@
An excellent taxonomy for types of constraints for CMDP can be found in [40]. A few key equations for the state-value function (3), state-action-value or Q function (4) and advantage function (5) are given below:
@@ 3.3 Lyapunov Functions
A continuously differentiable function $V:\mathcal{X}\mapsto\mathbb{R}$ is a Lyapunov function if [47]:
@@
where $x_{d}\in\mathcal{X}$ is the desired state for the system and $f_{cl}$ is the closed loop dynamics of the system. $\nabla V(x)f_{cl}(x)$ denotes the Lie derivative of $V$ along the function $f_{cl}$. Control Lyapunov functions provide formal guarantees for stabilizability of open-loop systems, which implies the existence of a controller that drives the closed loop system to stability. Formally, a control Lyapunov function (CLF) is a criterion that provides asymptotically stabilizing controllers for a general nonlinear dynamical system, which was first generalized by Sontag [48]. For a control affine form in Section 3.1, a CLF $V$ is a smooth (continuous and differentiable up to some order), proper ($V(x)\rightarrow\infty$ as $\|x\|\rightarrow\infty$) and positive definite ($V(x)>0$ and $V(0)=0$ for $x\neq 0$) function:

### v3

@@ Abstract
keywords:
Reinforcement Learning, Lyapunov Functions, Barrier Functions, Neural Networks (NNs),
@@ 1 Introduction
With increase in efficiency and function approximation capabilities of deep neural networks (DNN), reinforcement learning has seen increased research and some exciting developments over the last few decades [arulkumaran2017deep]. RL in its bare bones involves an agent iteratively interacting with an environment to compute a control or decision making policy by maximizing a reward function. RL has proven to be effective in computing control or decision policies for a wide variety of complex systems and environments such as robotics [ibarz2021train, wang2022distributed, abeyruwan2023sim2real, liu2022digital, muzio2022deep], computer vision [pinto2023tuning, le2022deep, tao2022evaluating], cyber-security [adawadkar2022cyber, piplai2022knowledge, tran2022cascaded, huang2022reinforcement, 10252253], energy management [ganesh2022review, liu2023energy, yang2022reinforcement, 9916919]), chess [bertram2022supervised, xu2022data, hammersborg2022reinforcement] and video games [taylor2014reinforcement, mnih2013playing, kaiser2019model] to name a few. However, when using RL there are several complications and theoretical gaps in terms of reproducibility, convergence guarantees, large amounts of data for training and large number of iterations required for convergence [henderson2018deep]. Even though RL is often known for its data-driven nature, it can be broadly classified into two approaches for deriving a control or decision policy namely, model-based (value and policy iteration) and model-free (value-function, policy search or gradient and actor-critic methods) approaches. Model-based approaches imply use of knowledge (partial or complete) or an approximation of the system to derive a control policy whereas, model-free approaches purely rely on data or experiences collected by interactions with the environment without using a model of the system [osinenko2022reinforcement]. It seems intuitive to have stronger guarantees for constraint satisfaction and computing a stabilizing policy while leveraging a model-based approach as the RL agent uses information from the system model. There has been some success in providing these guarantees for model-free approaches as well, notably in [cheng2019end]. A key challenge for policies derived through use of RL has been providing guarantees for stability of the system for a computed policy and satisfying system state constraints. Broadly speaking this challenge encapsulates the need for safety in RL for real-world applications and thus, encourages the research in this field.
@@
There are several methods in literature for computing a stabilizing policy and satisfying system constraints in RL. The simplest approaches offer a direct and intuitive method of filtering out “unsafe” actions and policies. The term “unsafe” can be referred to any action leading to constraint violation or system instability. In [curi2022safe], the authors propose a confidence safety filter for stochastic nonlinear systems. The key idea presented is to formulate safety constraints as cumulative costs which can then be expressed as cost constraints for RL. Backup policies in case of constraint violations are obtained for the safety filter by computing a safe policy through robust RL, which is then employed in a confidence-based safety filter. The authors in [vinod2022safe], use a convex optimization-based filtering to satisfy hard constraints for target tracking. They propose a two-level motion planner where a RL-based controller generates an input based on the target and quadratic programming (QP) based safety filter generates a safe input, regulation is performed on these two inputs produced. Similar to safety filter, supervisor-based safe RL methods often use human supervision [arakawa2018dqn, wu2023toward, luo2023human] to penalize or replace unsafe actions with desirable ones. These supervisor-based methods are often termed as human-in-loop RL. A formal approach to supervisor-based safe RL methods are classified as RL via shielding. Authors in [elsayed2021safe], propose a centralized (single shield for all agents in an environment) and factored (separate shields for a subsets of agents in an environment) shielding methodology for a multi-agent RL setting using linear temporal logic (LTL) for safety constraints. The scalability and learning performance are discussed for each methodology by the authors of mentioned work. In [alshiekh2018safe], authors express constraints as temporal logic and propose an algorithm for automatic synthesis of shields for the given temporal logic. This is further incorporated with RL and the results show increased learning performance compared to unshielded cases. However, a key drawback of this approach is that it relies on some prior information for model of the system to deem an action unsafe. Model predictive control (MPC) has long been effective in control theory for handling uncertainty in dynamical systems while satisfying operational constraints. There have been several notable works in combining the data-driven advantage of RL and the robustness of MPC, reviewed in [norouzi2023integrating]. RL using robust MPC is proposed in [zanon2020safe], where the authors use online data to compute safe design constraints to evaluate the cost function for RL, Q-function and value function are obtained via robust MPC. In [sawant2022bridging], combination of QP and MPC, one at a time with RL is explored. The policy and value functions of RL are approximated using MPC and QP instead of employing DNNs. Compared to lack of guarantees and explainability of DNNs, this approach provides structure to the approximations for policy and value functions.
@@
Control Lyapunov functions (CLF) [sontag1999control] and control barrier functions (CBF) [ames2019control] have proven to be an effective way to guarantee safety through stability of a closed-loop system and defining safe sets, respectively. Ensuring safe operation for dynamical systems involving human interaction, expensive equipment etc., requires strong guarantees during and after the learning process for RL agents. Using CBFs and CLFs to provide stability and safety guarantees is especially well suited for learning because both features may be stated using learnable functions. Considering their usefulness in a learning framework and their applicability to a wide variety of control frameworks, this paper reviews methods using Lyapunov and barrier functions to ensure safety and stability guarantees of a computed policy while using RL for decision-making problems. Furthermore, most real-world applications require safety during the training process (for example, navigation in unknown environments for drones, human presence in application etc.) which makes this problem crucial as it provides theoretical guarantees for safety. The key motivation for this review is to discuss methods employed in recent work, not covered in reviews and surveys mentioned in Section 2 and particularly focus on use of Lyapunov and barrier functions for RL to guarantee satisfying system constraints and stability in terms of the computed policy. We focus more on the theoretical approaches present in literature rather than the type of applications it is employed. To the best of author’s knowledge, such a review focusing on use of Lyapunov and barrier functions in safe RL does not exist and this research gap is addressed in this review. Rest of the review is organized as follows: Section 2 presents scope of work for different reviews and surveys considered. Section 3 discusses preliminary theory and notations briefly to provide the reader a background for further discussion. Section 4, 5 and 6 summarizes, classifies and discusses the current and prior work related to use of: (i) Lyapunov functions in RL; (ii) Barrier functions in RL; (iii) Lyapunov and barrier functions in RL, respectively. Section 7 provides a discussion on future research directions, current challenges and potential real-world applications. Finally, Section 8 presents conclusions to the review conducted.
@@ 2 Related Work
The literature in safe reinforcement learning has been fairly recent, spanning back to the last two decades. The earliest survey in this field by [garcia2015comprehensive], primarily focuses on safe RL by augmenting the optimization criterion and exploration process by including a risk parameter and state constraints. The survey also focuses on leveraging external knowledge in the form of expert policies and human intervention to guide the exploration process. The review by Liu et al. [liu2021policy] focuses on policy learning with constraints for model-free RL. The authors present a robust taxonomy for the types of constraints for a constrained Markov decision process (CMDP) and discusses various policy optimization approaches based on the type of constraints in CMDP. The authors broadly classifies constraints as cumulative and instantaneous depending on the long-term and short-term effects of constraints on the cost function, respectively.
@@
Brunke et al. [brunke2022safe] provides an extensive and one of the most comprehensive reviews covering safe learning for control from a robotics perspective, however, they do not cover key recent and prior works on combining Lyapunov and barrier functions in RL. To quote the authors, they provide a “bird’s eye view” of the field. Authors cover most of the literature associated with safe learning and provide excellent categorizations for levels of safety and various formulations for safe learning techniques. However, due to the vast range of topics considered, use of Lyapunov and barrier functions in RL is sparsely covered. The review does provide an excellent starting point for an overview into all possible directions for development in safe learning for control and RL. Gu et al. [gu2022review] focus on all developments in safe RL covering theory and applications. The authors formalize a “2H3W” problem, addressing the key problems and challenges pertaining to safe RL implementation. Convergence analysis and iteration complexity for various approaches to safe RL are explored for model-free and model-based RL along with applications and available benchmarks. The authors also discuss application of safe RL in multi-agent settings, discussing possible research avenues. The review is detailed, however, it covers only a few approaches pertaining to use of Lyapunov and barrier functions.
@@
The authors in [osinenko2022reinforcement], present an overview for RL with guarantees. The authors cover three approaches for stability of the system for a computed policy and constraint satisfaction namely, RL with supervisors, model predictive control (MPC)-based RL and Lyapunov-based RL. The review presents few instances of using CBF with CLF, but it primarily gives an overview of key challenges in consideration of model-based and model-free RL while providing safety guarantees. Anand et al. [anand2021safe] presents a concise and focused review on safe learning for control using CLF and CBF. The authors cover safety in RL, online supervised learning (SL) and offline SL with incorporation of CLF, CBF and combining CLF with CBF. However, the review presents a few notable approaches for combining CLF and CBF with RL and only covers one CMDP based formulation. The excellent survey by Dawson et al. [dawson2022safe] covers a great depth of literature available on learning-based control using CLF, CBF and contraction methods. The review provides excellent theoretical overview of different formulations and implementations using control certificates but is more focused towards learning-based control (adaptive controls) from a control theoretic perspective compared to RL.
@@ 3.1 Definition of Safety
A general continuous time nonlinear dynamical system is represented as $\dot{x}=f(x,u)$, where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is state of the system, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control input and $f:\mathcal{X}\times\mathcal{U}\mapsto\mathcal{X}$ is the state transition function (assumption: Locally Lipschitz in $x$ and $u$) [dawson2022safe]. The control affine form of nonlinear dynamical system can represented as $\dot{x}=f(x)+g(x)u$, where $g:\mathbb{R}^{n}\mapsto\mathbb{R}^{n\times m}$ (assumption: Locally Lipschitz in $x$ and $u$).
@@ 3.2 Constrained Markov Decision Process (CMDP) and Safe RL
A Markov Decision Process (MDP) can be denoted by the tuple $\mathcal{S},\mathcal{A},\mathcal{R},\mathbb{P},\mu,\gamma>$  [meyn2022control], where $\mathcal{S}$ and $\mathcal{A}$ denote the set of states and actions, respectively. $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ denotes the reward function, $\mathbb{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ denotes the transition probability function, $\mu:\mathcal{S}\mapsto[0,1]$ is the initial probability distribution and $\gamma$ denotes the discount factor for future rewards. A policy $\pi:\mathcal{S}\mapsto\mathcal{P}(A)$ is a mapping from states to a probability distribution over actions and $\pi(a_{t}|s_{t})$ is the probability of taking action $a$ under state $s$ at time $t$. Reinforcement learning (RL) involves an agent successively interacting with its environment, to accomplish the aim for learning a policy to maximize a given reward criteria. The goal of RL is to learn a policy $\pi$ that maximizes the discounted cumulative reward:
@@
where $\tau$ denotes a state-action trajectory, and $\tau\sim\pi$ denotes trajectories sampled from the policy $\pi$. A CMDP is an extension of MDP with an additional constraint set of cost functions $\mathcal{C}$ [altman2021constrained]. The augmented tuple for a CMDP is then denoted by $\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{C},\mathbb{P},\mu,\gamma>$. The cost functions $\mathcal{C}_{i}\in\mathcal{C}$, $\mathcal{C}_{i}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ are constraint associated cost functions. We call an action $a$ feasible if $a\in\mathcal{A}$ satisfies all of its constraints. The objective for safe RL is then to find a policy to maximize the long-term rewards while satisfying the constraints, denoted as:
@@
An excellent taxonomy for types of constraints for CMDP can be found in [liu2021policy]. A few key equations for the state-value function (3), state-action-value or Q function (4) and advantage function (5) are given below:
@@ 3.3 Lyapunov Functions
A continuously differentiable function $V:\mathcal{X}\mapsto\mathbb{R}$ is a Lyapunov function if [khalil2002nonlinear]:
@@
where $x_{d}\in\mathcal{X}$ is the desired state for the system and $f_{cl}$ is the closed loop dynamics of the system. $\nabla V(x)f_{cl}(x)$ denotes the Lie derivative of $V$ along the function $f_{cl}$. Control Lyapunov functions provide formal guarantees for stabilizability of open-loop systems, which implies the existence of a controller that drives the closed loop system to stability. Formally, a control Lyapunov function (CLF) is a criterion that provides asymptotically stabilizing controllers for a general nonlinear dynamical system, which was first generalized by Sontag [sontag1989universal]. For a control affine form in Section 3.1, a CLF $V$ is a smooth (continuous and differentiable up to some order), proper ($V(x)\rightarrow\infty$ as $\|x\|\rightarrow\infty$) and positive definite ($V(x)>0$ and $V(0)=0$ for $x\neq 0$) function: